OpenStudy (anonymous):

“Simplify the expression: 310 – 20 ÷ 2 × 5 + 13” Below is a student answer: 310 – 20 ÷ 2 × 5 + 13 310 – 20 ÷ 10 + 13 310 – 2 + 13 310 – 15 295 Part A) The student made two mistakes. EXPLAIN what those two mistakes were in regards to the order of operations. Part B) Simplify the expression yourself using the order of operations.

OpenStudy (ipwnbunnies):

Do you know about order of operations, and PEMDAS?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Should they have started from the back first?

OpenStudy (anonymous):

Or the front?

OpenStudy (ipwnbunnies):

PEMDAS: Parentheses, Exponents, Multiplication, Division, Addition, Subtraction. It's an order in which you're supposed to solve math problems, like this. Also, M & D, and A & S can be switched up if one of them shows up first in the problem.

OpenStudy (ipwnbunnies):

In the original problem, you have 20/2*5, which is where the first mistake is. The student did the 2*5 part first, to get 20/10. You cannot do this, according to order of operations. You must do the 20/2 division first, because it comes up first in the problem.

OpenStudy (ipwnbunnies):

Instead of 310 – 20 ÷ 10 + 13, It should be 310 - 10 * 5 + 13

OpenStudy (anonymous):

Thank you

OpenStudy (ipwnbunnies):

The second problem comes during the addition and subtraction part. You have to do it in order as it shows up in the problem. The student did the addition on the right side first, but you have to go from left to right.

OpenStudy (anonymous):

So when i Simplify the expression it should be: 310 – 20 ÷ 2 × 5 + 13” 310 - 20 / 10 plus 13

OpenStudy (ipwnbunnies):

Noooo, that was the problem we found. You would have to do 20/2 first, because it comes up first.

OpenStudy (anonymous):

oh so you do the one right next to the first number?

OpenStudy (ipwnbunnies):

Yeah, after you do the 20/2, you'll multiply that answer by 5. Since divison and multiplication operations come before addition and subtraction, I showed you in PEMDAS.

OpenStudy (anonymous):

i get it now, you want me to try again?

OpenStudy (ipwnbunnies):

Yeah, lemme see if you can evaluate the problem correctly, which is what you're supposed to do for Part B.

OpenStudy (anonymous):

310 – 20 ÷ 2 × 5 + 13 310 - 10 * 5 plus 13 310 - 50 plus 13 310 - 50 lol it got hard

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

310 - 63

OpenStudy (ipwnbunnies):

Noooo, you can't do that either. That was the problem the student made. You have to go from left to right. So you have to do 310 - 50 first.

OpenStudy (anonymous):

can u copy the right part of it and paste it

OpenStudy (ipwnbunnies):

310 - 50 + 13 The last part of PEMDAS is Addition and Subtraction. But you still have to do it in order of the problem. 310 - 50 = 260 260 + 13 = 273

OpenStudy (anonymous):

i quittt sorry

OpenStudy (anonymous):

310 – 20 ÷ 2 × 5 + 13 310 - 10 * 5 plus 13 310 - 50 = 260 260 + 13 = 273 ???

OpenStudy (anonymous):

or 310 – 20 ÷ 2 × 5 + 13 20 / 2 = 10 10 * 5 = 50 310 - 50 = 260 260 + 13 = 273

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

They're both right. I was just breaking it up so you can see each individual section I was doing.

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

what one should i put for the test